A typical capacitor in a memory cell have a capacitance of 3 x 10-14 F. If the voltage across the capacitor is 0.5 V, what is the number of electrons that must move to the capacitor to charge it? 


1
Expert's answer
2022-01-21T13:48:42-0500

Given:

"C=3\times 10^{-14}\: \rm F"

"V=0.5\: \rm V"

"e=1.6\times 10^{-19}\:\rm C"


The charge on the capacitor plates

"q=CV=3\times 10^{-14}\times 0.5=1.5\times 10^{-14}\:\rm C"

The number of electrons

"N=\frac{q}{e}=\frac{1.5\times 10^{-14}\:\rm C}{1.6\times 10^{-19}\:\rm C}=93750"